8. A person is hearing two sound waves simultaneously One has a period of 1.50 ms and the other one a period of 1.54 ms. What is the period of the beat due to these two waves?
Answer #1

tregue T, = 1-50ms = 1.50 X 16-3 sec frequency (f) = + fi = t = -50 x logo Hz T2 = 1.54ms - 1.54 x103 sec f2 = F = 1.56X163 =